OpenStudy (anonymous):

simplify ((2x)/(x^2-6+9))-(1/(x+1))-(8/(x^2-2x-3))

OpenStudy (anonymous):

\[\frac{ 2x }{ x^2-6x+9 } - \frac{ 1 }{ x+1 } - \frac{ 8 }{ x2-2x-3 }\]